The pitch of a roof is its slope. Find the pitch of 12/15
Answer by math_helper(2461)

Roof pitch is almost always expressed as inches per foot (so some # over 12-inches).
Given 12/15, we need to get it in terms of x /12:
12/15 = x/12
144/15 = x
9.6 = x
So the roof pitch is "9.6 inches per foot" or 9.6 / 12